OpenStudy (anonymous):

Write the first trig function in terms of the second. sinx;cosx I don't understand the steps to get a final answer.

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sin^2(x) + \cos^2(x) = 1\]

jimthompson5910 (jim_thompson5910):

do you see how to isolate sin(x)?

OpenStudy (anonymous):

Do I subtract cos?

jimthompson5910 (jim_thompson5910):

good first step, then what

jimthompson5910 (jim_thompson5910):

doing that will give you \[\Large \sin^2(x) + \cos^2(x) = 1\] \[\Large \sin^2(x) = 1 - \cos^2(x)\] what do we need to do next to isolate sin(x)?

OpenStudy (anonymous):

square root both sides?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

\[\Large \sin^2(x) + \cos^2(x) = 1\] \[\Large \sin^2(x) = 1 - \cos^2(x)\] \[\Large \sin(x) = \sqrt{1 - \cos^2(x)}\]

OpenStudy (anonymous):

so that would be the final answer?

jimthompson5910 (jim_thompson5910):

yes it would